3 square root sign 64x^7 divided by 3 square root sign 4x
Answer by checkley77(12844)

[HOPE THESE EXPLANATIONS HELP.]
3SQRT(64X^7)/3SQRT(4X) CANCEL OUT THE 3's in the numerator & denominator.
Factor all squares under the sqrts. (64=8*8, x^7=x^3*x^3*x^1,4=2*2)
SQRT(8*8*X^3*X^3*X)/SQRT(2*2*X)
Extract the square roots from under the square root sign. (8, x^3 & 2)
8X^3SQRTX/2SQRTX
DIVIDE THE NUMERATOR & DENOMINATOR BY 2 & CANCEL OUT THE SQRTX TERMS.
4X^3 ANSWER.